# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v2

@@ II-B Belief Propagation (BP) Decoding
    At each iteration, an error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$ is made according to the hard decision on ${\mathbf{q}}_{i}$. If a valid error estimate is generated, it will be accepted as the solution; otherwise, go to the next iteration starting from step 2).
Steps 2) and 3) will be iterated for a maximum number of ${\mathrm{T}}$ iterations, where ${\mathrm{T}}$ is chosen in advance. If no valid estimate is obtained after ${\mathrm{T}}$ iterations, BP will claim a failure.
III Order-statistic decoding based on quaternary belief propagation
In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{H}}$ and an error syndrome ${\mathbf{s}}$ with $2n$ binary variables in ${\mathbf{e}}$. The linear system has $n-k$ independent constraints, impliying that there are $n+k$ degrees of freedom in the vector ${\mathbf{e}}$. Given the values of these $n+k$ bits, one can reconstruct ${\mathbf{e}}$. If BP fails to provide a valid error estimate after ${\mathrm{T}}$ iterations, OSD will be utilized.
In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits. Then, we can use these coordinates to generate a valid error that matches the syndrome. To accomplish this, it is necessary to establish the concept of reliability order as a means of prioritizing the coordinates based on their reliability.
III-A Reliability order based on BP4
To utilize the quaternary distribution generated by BP4 for each qubit error in OSD, we propose a method that employs the hard-decision history from all BP iterations and the output probabilities from the last BP iteration.
Suppose that BP fails to provide a valid error after a maximum of ${\mathrm{T}}$ iterations. At this point, we have the output distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$, $i=1,2,\dots,n$ from the last BP iteration. Additionally, we have the hard-decision results for all ${\mathrm{T}}$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,{\mathrm{T}}$. For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in [1, 43], we define two types of measures in the following.
@@ III-A Reliability order based on BP4
In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$. For example, if the hard-decision outputs at the first qubit over the last five iterations are $X,Y,Z,Z,Z$, then the final hard decision output is $Z$ and $\bm{\eta}_{1}=3$. This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops.
@@
Note that $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$ is a binary distribution indicating the likelihood of an $X$ error occuring at the $i$-th qubit and similarly for $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$.
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, error rate $\epsilon$, maximum number of iterations ${\mathrm{T}}$.
Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$ and hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$.
@@
Choose initial error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for $i=1,\dots,n$. Let $\bm{\eta}_{i}=1$ for $i=1,\dots,n$. Let ${\mathbf{h}}_{i}=I$ for $i=1,\dots,n$. Let ${\mathbf{q}}_{i}=\mathbf{0}$ for $i=1,\dots,n$.
@@
Definition 4.
For $1\leq i,j\leq n$ and $a,b\in\{X,Z\}$, error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{e}}_{j}^{b}$ if $\bm{\eta}_{i}>\bm{\eta}_{j}$, or if $\phi^{a}(i)\geq\phi^{b}(j)$ when $\bm{\eta}_{i}=\bm{\eta}_{j}$.
This definition means that the hard-decision history reliability is used to rank the error coordinates, with the soft reliability serving as a tie-breaker.
Let $\tilde{{\mathbf{e}}}$ be the hard-decision output from BP4 after ${\mathrm{T}}$ iterations. If $\tilde{{\mathbf{e}}}$ does not match the syndrome, OSD will be applied. We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is referred to as OSD4-00.
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector $\tilde{{\mathbf{e}}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$.
@@
1:Sort the error bits in ascending order of reliability. Construct a column permutation function $\pi$ corresponding to the sorting order. Calculate $\pi(\tilde{{\mathbf{e}}})$ and $\pi({\mathbf{H}}\Lambda)$.
2:Perform Gaussian elimination on $\pi({\mathbf{H}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{H}}\Lambda))$, where $R$ represents the corresponding row operations in the Gaussian elimination. If necessary, apply a column permutation function $\mu$ to ensure that the first $n-k$ columns of the resulting matrix are linearly independent. Thus we have
$\mu(R(\pi({\mathbf{H}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (7)
for some ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$, where the last $m-(n-k)$ rows are all zeros.
3:Let ${\mathbf{s}}^{\prime}=R({\mathbf{s}})$ so that the parity-check condition (2) becomes
${\mathbf{s}}^{\prime}=\mu(R(\pi({\mathbf{H}}\Lambda)))(\mu(\pi({\mathbf{e}})))^{\top}.$ (8)
Remove the last $m-(n-k)$ entries of ${\mathbf{s}}^{\prime}$.
4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~{}~{}\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits. Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 8 to select feasible unreliable bits
@@
We also define a reliability order that accounts for only the soft reliability functions of BP4.
Definition 5.
Error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{e}}_{j}^{b}$ in the marginal distribution if $\phi^{a}(i)\geq\phi^{b}(j)$.
The OSD based on Definition 5 is referred to as mOSD4.
Other alternative metrics can also be considered, such as
[…]
The reliability of qubit $i$ is assessed based on the quaternary entropy of its belief distribution $\mathbf{q}_{i}$ (Entropy) or based on the maximum probability of the distribution (Max). For our purpose, the hard-decision output of BP is first sorted in the Pauli basis based on the reliability metric (either Entropy or Max) and then transformed into a binary vector for OSD [43]. We will compare these reliability metrics in Section VI-D.
III-B OSD4-$w$

### v3

@@ III-A Density-Evolution-Guided Reliability Heuristic
    Variable update: Each variable node $v_{j}$ updates its estimate according to the unanimous recommendations of its neighboring checks:
    • $X$-flip: if all incoming messages from $Z$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}X$;
    • $Z$-flip: if all incoming messages from $X$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}Z$.
    (If both updates occur, a $Y$ correction is effectively applied up to a global phase.)
  4. 4.
    Message update: Set $m_{v_{j}\to c_{i}}^{(\ell+1)}=1$ if $\hat{E}_{j}$ anticommutes with the stabilizer of $c_{i}$, and 0 otherwise.
  5. 5.
    Termination: The algorithm stops if the current estimate $\hat{\mathbf{E}}$ is valid.
For a CSS-type stabilizer code, $X$-type and $Z$-type errors are processed independently by the modified Gallager-A algorithm. The correlations between $X$ and $Z$ errors induced by the depolarizing channel are therefore neglected at this stage. Accordingly, we focus on the analysis of $X$-error decoding, which reduces to a classical binary decoding problem: estimating an $n$-bit vector ${\mathbf{e}}=(e_{1},\dots,e_{n})\in\{0,1\}^{n}$ from its syndrome.
Suppose that the $Z$-type stabilizers define a $(d_{c},d_{v})$-regular LDPC code, where each check node is connected to $d_{c}$ variable nodes and each variable node is connected to $d_{v}$ check nodes.
Proposition 2.
[Heuristic Density Evolution Analysis] Consider a quantum code of length $n$ under depolarizing noise with rate $\epsilon$. Assume that the Tanner graph induced by the $Z$ stabilizers is locally tree-like and that decoding is performed using the modified Gallager-A algorithm. Let $x^{(\ell)}$ denote the probability that a variable-to-check message is incorrect at iteration $\ell$, and let $y^{(\ell)}$ denote the probability that a check-to-variable message is incorrect. Then the message error probabilities evolve according to
$y^{(\ell)}$ $=\frac{1-(1-2x^{(\ell)})^{d_{c}-1}}{2},$ (6)
$x^{(\ell+1)}$ $=p_{a}\big[1-(1-y^{(\ell)})^{d_{v}-1}\big]+(1-p_{a})(y^{(\ell)})^{d_{v}-1},$ (7)
where $p_{a}=\frac{2\epsilon}{3}$ is the probability that a Pauli error anticommutes with a $Z$-type stabilizer, and $x^{(0)}=p_{a}$.
The recursion (6)–(7) follows from the standard density-evolution analysis of Gallager-A decoding for regular LDPC codes over the binary symmetric channel [15, 58, 59].
At error rates below the heuristic threshold, density evolution predicts that most variable nodes converge rapidly, leaving a fraction proportional to the channel error probability $\epsilon$ unresolved. The expected number of erroneous variable nodes at iteration $\ell$ is $n\cdot x^{(\ell)}$, and after an average convergence time $\ell_{\mathrm{succ}}$ the residual error set has size roughly $n\cdot x^{(\ell_{\mathrm{succ}})}$. For finite-length codes, cycles in the Tanner graph cause deviations from the infinite-length density evolution prediction. As $\epsilon$ decreases, $x^{(\ell_{\mathrm{succ}})}$ rapidly vanishes, leaving extremely sparse residual errors.
Unlike classical LDPC ensembles, stabilizer Tanner graphs inherently contain many short cycles due to commutation constraints. As a result, the above density-evolution analysis should be interpreted as a heuristic guideline rather than a rigorous asymptotic prediction. Although one could in principle extend DE to full belief propagation with real-valued LLRs, the Gallager-A approximation already captures the main qualitative behavior of the decoding process. Nonetheless, our heuristic results serve as a useful reference for BP algorithms that exploit memory effects to mitigate the impact of short cycles [32, 77].
III-B Reliable subset reduction for the syndrome decoding problem
In the syndrome decoding problem, we seek a binary vector ${\mathbf{e}}\in\{0,1\}^{2n}$ satisfying
${\mathbf{H}}\Lambda{\mathbf{e}}^{\top}={\mathbf{s}},$
where $\mathbf{H}$} has rank nkn-k and hence the solution space has dimension $n+k$. When BP fails to converge to a valid solution within $\mathrm{T}$} iterations, a postprocessing algorithm such as OSD is typically required.
Motivated by density evolution, we expect that after a few BP iterations, most variables have already converged to highly reliable values, while only a small subset remains ambiguous. We denote by $R\subset\{e_{1},\dots,e_{2n}\}$ a subset of reliable variables, whose values are assumed to be correct and fixed. The remaining variables are treated as unknowns.
By eliminating the reliable variables from the parity-check equations, the original syndrome decoding problem can be reduced to a much smaller linear system. We refer to this procedure as reliable subset reduction (RSR). The number of remaining variables after reduction is called the effective length.
Algorithm 1 RSR
Input: parity-check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome $\mathbf{s}$}, reliable subset $R$.
Output: reduced matrix $\tilde{{\mathbf{H}}}$, reduced syndrome $\tilde{{\mathbf{s}}}$, and permutation $\sigma$.
Steps:
1:Let $\sigma$ be a column permutation such that $\sigma({\mathbf{e}})=\begin{bmatrix}{\mathbf{e}}^{\prime}&{\mathbf{e}}^{R}\end{bmatrix}$, where ${\mathbf{e}}^{R}$ is the reliable subset of $v$ bits, and $\mathbf{e}$$^{\prime}$ is the remaining 2nv2n-v bits. Let $\sigma(\hat{{\mathbf{e}}})=\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{R}\end{bmatrix}$ be a permuted error estimate, which contains the reliable bits in ${\mathbf{e}}^{R}$.
2:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{H}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{H}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$. The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the reliable bits. Assume that $\tau({\mathbf{s}})=\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{R}\end{bmatrix}$, where ${\mathbf{s}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{s}}^{R}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
3:Verify the parity check condition related to the highly reliable bits.
if ${\mathbf{s}}^{R}\neq{\mathbf{C}}\big({\mathbf{e}}^{R}\big)^{\top}$ then return “Failure";
else we have a new parity-check condition
  where $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big({\mathbf{e}}^{R}\big)^{\top}$.
  if (8) is solvable then return $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, and $\sigma$;
  else return “Failure".
  end if
end if
Equation (8) follows directly from the block decomposition
$\tau(\sigma({\mathbf{H}}))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\end{bmatrix}$,
which separates the contributions of the unknown variables $\mathbf{e}$$^{\prime}$ and the fixed reliable variables ${\mathbf{e}}^{R}$.
There are two possible failure modes in RSR:
  1. (i)
    Stage-1 failure: the reliable bits are inconsistent with the syndrome, i.e., ${\mathbf{s}}^{R}\neq{\mathbf{C}}({\mathbf{e}}^{R})^{\top}$.
  2. (ii)
    Stage-2 failure: the reduced linear system $\tilde{{\mathbf{H}}}(\hat{{\mathbf{e}}}^{\prime})^{\top}=\tilde{{\mathbf{s}}}$ is rank-deficient or otherwise unsolvable.
Both conditions are explicitly verified in Step 3 of Algorithm 1. The stage-1 failure corresponds to checking the sparse consistency condition
${\mathbf{s}}^{R}={\mathbf{C}}({\mathbf{e}}^{R})^{\top},$
which only requires a sparse matrix–vector multiplication and can be performed efficiently. The stage-2 failure concerns the rank and pivot structure of $\tilde{{\mathbf{H}}}$. In many applications, this information must be extracted in order to solve the reduced linear system, for example through Gaussian elimination or related procedures. Therefore, this check can often be performed with little or no additional overhead when RSR is used together with a downstream decoder that requires such a step (e.g., OSD [44], LSD [55]).
If RSR succeeds, the reduced system
$\tilde{{\mathbf{H}}}(\hat{{\mathbf{e}}}^{\prime})^{\top}=\tilde{{\mathbf{s}}}$
is passed to a downstream decoder to obtain an estimate $\hat{{\mathbf{e}}}^{\prime}$. The full error estimate is recovered as
$\hat{{\mathbf{e}}}=\sigma^{-1}\big(\hat{{\mathbf{e}}}^{\prime},{\mathbf{e}}^{R}\big).$
III-C Reliability order based on BP4
To quantify the reliability of each error-variable estimate, we combine two sources of information: (i) the hard-decision trajectory of BP across all iterations, and (ii) the belief distribution produced at the final BP iteration.
Suppose that BP terminates without finding a valid solution after a maximum of $\mathrm{T}$} iterations. At this point, BP provides a belief distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ for each qubit $i=1,\dots,n$, as well as a sequence of hard decisions ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $j=1,\dots,{\mathrm{T}}$. Here ${\mathbf{h}}_{i}^{(j)}$ denotes the Pauli error assigned to qubit $i$ at iteration $j$.
Our goal is to assign a reliability score separately for the $X$ and $Z$ components of each qubit. Following [1, 44], we introduce two complementary measures.
@@ III-C Reliability order based on BP4
In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$. For example, if the hard-decision outputs at the first qubit over the last five iterations are $X,Y,Z,Z,Z$, then the final hard decision output is $Z$ and $\bm{\eta}_{1}=3$. This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops. A larger $\eta_{i}$ indicates that the estimate at qubit $i$ has stabilized over a longer period and is therefore more reliable.
@@
Here $\{q_{i}^{X}+q_{i}^{Y},\;q_{i}^{I}+q_{i}^{Z}\}$ forms a binary distribution corresponding to whether an $X$ error is present on qubit $i$, and similarly $\{q_{i}^{Z}+q_{i}^{Y},\;q_{i}^{I}+q_{i}^{X}\}$ corresponds to a $Z$ error.
Definition 5.
An error bit ${\mathbf{e}}_{i}^{a}$, where $i\in\{1,\dots,n\}$ and $a\in\{X,Z\}$, is said to be highly reliable if its hard-decision reliability is either $\mathrm{T}$} or ${\mathrm{T}}+1$, and its soft reliability $\phi^{a}(i)$ satisfies
$\phi^{a}(i)\geq\theta,$
where $\theta\in(0,1)$ is a predefined reliability threshold.
An error variable has hard reliability $\mathrm{T}$} if its hard decision remains constant throughout all BP iterations. A value of ${\mathrm{T}}+1$ occurs when the variable stays equal to the identity operator $I$ from initialization through all iterations. The threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate is not only stable in its hard-decision history but also strongly supported by the final soft information.
Definition 6.
For $1\leq i,j\leq n$ and $a,b\in\{X,Z\}$, the error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than ${\mathbf{e}}_{j}^{b}$ if either
$\eta_{i}>\eta_{j},$
or
$\eta_{i}=\eta_{j}\quad\text{and}\quad\phi^{a}(i)\geq\phi^{b}(j).$
That is, the stability of the hard-decision trajectory at the qubit level provides the primary ranking, while the final soft reliability acts as a tie-breaker between individual $X$ and $Z$ components.
@@
There exist various reliability metrics based on the (weighted) accumulation of log-likelihood ratio (LLR) signs across BP iterations [78, 79, 80, 81]. Our hard-decision reliability can be viewed as a natural extension of this idea, focusing on the length of the most recent run in the hard-decision sequence. The reliability order in Definition 6 can also be adapted to incorporate alternative metrics, such as Entropy or Max [44]. In these variants, the reliability of qubit $i$ is quantified either by the quaternary entropy of its belief distribution $\mathbf{q}_{i}$ (Entropy) or by the maximum probability within the distribution (Max).
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, maximum number of iterations $\mathrm{T}$}, and initial error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for $i=1,\dots,n$.
[…]
@@
Let $\bm{\eta}_{i}=1$ for $i=1,\dots,n$. Let ${\mathbf{h}}_{i}=I$ for $i=1,\dots,n$ be the initial hard decisions. Let ${\mathbf{q}}_{i}=\mathbf{0}$ for $i=1,\dots,n$.
@@
IV Order-statistic decoding based on quaternary belief propagation
Once BP fails to produce a valid error estimate, combinatorial post-processing techniques are required. These techniques operate on the residual errors left by BP and can include methods such as OSD, LSD, or other syndrome-based decoders. In this section, we first review the basic OSD procedure, which relies on the reliability order defined in Section III-C. When combined with the RSR method introduced in Section III-B, OSD can operate on a reduced system of variables, thereby lowering computational complexity.
In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits. Then, we can use these coordinates to generate a valid error that matches the syndrome. To accomplish this, we use the reliability orders defined in Section III-C.
Let $\tilde{{\mathbf{e}}}$ be the hard-decision output from BP4 after $\mathrm{T}$} iterations. If $\tilde{{\mathbf{e}}}$ does not match the syndrome, OSD will be applied. We now present our OSD algorithm based on the reliability orders as in Algorithm 3, which is referred to as OSD4-0.
@@ IV Order-statistic decoding based on quaternary belief propagation
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank nkn-k, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector $\tilde{{\mathbf{e}}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
[…]
@@
[…]
2:Perform Gaussian elimination on $\pi({\mathbf{H}}\Lambda)$ and let the output be denoted by $\rho(\pi({\mathbf{H}}\Lambda))$, where $\rho$ represents the corresponding row operations in the Gaussian elimination. If necessary, apply a column permutation function $\mu$ to ensure that the first nkn-k columns of the resulting matrix are linearly independent. Thus we have
$\mu(\rho(\pi({\mathbf{H}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (11)
for some ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$, where the last m(nk)m-(n-k) rows are all zeros.
3:Let ${\mathbf{s}}^{\prime}=\rho({\mathbf{s}})$ so that the parity-check condition (2) becomes
${\mathbf{s}}^{\prime}=\mu(\rho(\pi({\mathbf{H}}\Lambda)))(\mu(\pi({\mathbf{e}})))^{\top}.$ (12)
Remove the last m(nk)m-(n-k) entries of $\mathbf{s}$$^{\prime}$.
4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~~\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of nkn-k bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits. Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 12 to select feasible unreliable bits
@@
IV-A OSD4-$w$